OpenStudy (anonymous):

Solving Equations and Inequalties. Please help. 5^x= 4^(x+3)

OpenStudy (anonymous):

use logs... \[\LARGE \log_{5}(5^x)=\log_{5}(4)^{x+3}\] \[\LARGE x\cdot 1=(x+3)\cdot \log_5(4) \] \[\LARGE \frac{x}{x+3}=\log_5(4)\] I might be going wrong... I'll leave someone else here :(

OpenStudy (anonymous):

thank you a bunch thats it :)

OpenStudy (anonymous):

i got another one is that cool?

OpenStudy (anonymous):

don't you have to solve x= something even if it's not integers...

OpenStudy (anonymous):

ya but it works ok. it worked right for me.

OpenStudy (anonymous):

the only problem is that you have to use calculator, and unfortunately I can't get in it right now.. seems to have a little trouble. As soon as I can use it I'll leave you a solution ;)

OpenStudy (anonymous):

o ok. lol i can acces mine how would i put in as the whole thing or side to side or wat?

OpenStudy (anonymous):

Only \[\LARGE \log_54=\] I need to know this one ... :(

OpenStudy (anonymous):

Never mind... I got it... 0.8613531161 Now I'll do it lol

OpenStudy (anonymous):

\[\LARGE \frac{x}{x+3}\approx 0.86\] \[\LARGE 0.86x+2.58\approx x\] \[\LARGE 0.86x-x\approx -2.58\] \[\LARGE -0.14x\approx -2.58\] \[\LARGE 0.14x\approx 2.58\] \[\LARGE x\approx \frac{2.58}{0.14 }\] \[\LARGE x\approx 18.42\] I guess that's an UGLY answer lol

OpenStudy (anonymous):

lol its better than nothing. lol i got another one but it says solve by graphing with a calculator and round decimal to 3 decimal places. 9^(2x)-2=8^x

OpenStudy (anonymous):

this is why I called it UGLY answer... anyway integer number does it fit lol \[\Large \begin{array}{l}{5^x} = {4^{x + 3}}\\{5^{18.42}} = {4^{18.42 + 3}}\\7.4994166019313 \cdot {10^{12}} = {4^{21.42}}\\7.4994166019313 \cdot {10^{12}} = 7.8727237810908 \cdot {10^{12}}\end{array}\]

OpenStudy (anonymous):

still \[\LARGE 7.49\approx 7.87\] :P

OpenStudy (anonymous):

lol nice